OpenStudy (anonymous):

The volume of a cylinder is V = pirsquaredh. Find h if pi=3.14, V = 128, and r = 8. Round to the nearest hundredth. A. 0.56 B. 0.64 C. 0.58 D. 0.52

OpenStudy (anonymous):

B

OpenStudy (anonymous):

The volume of a cylinder isn't V = pi^2h

OpenStudy (anonymous):

V = (pi)(r^2)(h)

OpenStudy (anonymous):

so, plug in the numbers 128 = pi(64)(h)

OpenStudy (anonymous):

and solve for h

OpenStudy (anonymous):

\[V=\pi*r^2*h\] So you would substitue the values in: V = 3.14(64)*h 128=200.96*h You get 0.6369426751592357 And rounding to the nearest tenth gives you: 0.64 Which is B.

OpenStudy (anonymous):

Thanks You Guys :)